The ratio of robins to blackbirds that visit our feeder is 4 to 7. If there are a total of 132 birds, how many robins visit the feeder?
Answer by ptaylor(2198)

Let x=number of robins that visit
Now we know that out of every 11 birds, there are 4 robins, so:
4 is to 11 as x is to 132 and this is written as follows:
4/11=x/132 cross multiply (or multiply each side by 11*132)
4*132=11x
528=11x divide each side by 11
x=48 --------------number of robins that visit
CK
4/11=48/132
4/11=4/11
Hope this helps---ptaylor